Two identical particles each of mass M and charge Q are placed some distance apart. If they are in equilibrium under mutual gravitational and electric force then calculate the order of Q/M in SI system

Grade:12th pass

2 Answers

Arun
4 years ago
electric force depend upon the amount of charge not mass
let both particle are r distance apart
electric force is equal to 
=kQQ/r^2
gravitational force is equal to
=GMM/r^2
if both are equilibrium then
kQQ/r^2=GMM/r^2
Q^2/M^2=G/k
here G=6.67x10^-11
k=9x10^9 Nm^2/C^2
so G/k=6x10^-11/9x10^9
2/3x10^-20
Q/M=(2/3)^1/2 x10^-10
 
 
Khimraj
4 years ago
r
epulsive electrostatic force:    F1=kQ^2/r^2                 
       attractive gravitational force: F2=Gm^2/r^2
                                            
 
for equilibrium :F1=F2
kQ2/r2=Gm2/r2
=> Q/m = (6.6*10^-11/9*10^9)1/2
=> Q/m=0.8*10-10
Hence, ORDER OF Q/m=10-10